# Segmented Trajectory Optimization for Autonomous Parking in Unstructured Environments

### v1

@@ II-A Vehicle Geometry and Dynamics
Figure 1: Vehicle geometry and state. $P$: rear-axle center; $O^{\prime}$: instantaneous center; $FL$: front-left corner; $FR$: front-right corner; $RL$: rear-left corner; $RR$: rear-right corner; $L_{f}$: distance from rear axle to front bumper; $L_{r}$: distance from rear axle to rear bumper; $W$: vehicle width.
@@ II-D3 Switching Points
where $i$ ranges from 1 to $M-1$.
@@ II-D4 Reference Trajectory Proximity
$\begin{cases}0\leq v_{k}^{(i)}\leq v_{\text{max}},&\text{if $i$-th segment is forward}\\ v_{\text{min}}\leq v_{k}^{(i)}\leq 0,&\text{if $i$-th segment is backward}\end{cases}$ (9)
@@ II-D5 Collision Avoidance
A safe convex corridor is constructed around the reference trajectory using convex polygons. Let $P_{k}^{(i)}$ be the polygon enclosing $\mathbf{x}_{r,k}^{(i)}$, expressed as $P_{k}^{(i)}=\big{\{}\mathbf{p}\mid A_{k}^{(i)}\mathbf{p}\leq b_{k}^{(i)}\big{\}}$, where $A_{k}^{(i)}\in\mathbb{R}^{n\times 2}$ and $b_{k}^{(i)}\in\mathbb{R}^{n}$ are known matrices, and ${n}$ is the number of enclosing hyperplanes at this point. The method to construct the safe convex corridor will be presented in Section III-B. Collision avoidance is enforced by constraining all vehicle corners to remain within the polygon:
[…]
@@ 3 $\bm{X}_{r}\leftarrow$ PlanSimpleSpeed$(\rho)$
5       $\bm{P}\leftarrow$ ConstructConvexCorridor$(\bm{O}_{b},\bm{X}_{r})$
6       $\bm{X},\bm{U}\leftarrow$ SolveApproximateQP$(\bm{P},\bm{X}_{r})$
7       $e\leftarrow$ EvalErrorWithRK4$(\bm{X},\bm{U})$
@@ 11 $P\leftarrow\mathbb{R}^{2}$
13       $\mathbf{p}_{*}\leftarrow$ FindClosestPoint$(\bm{O}_{b,\text{remain}})$
14       $\varepsilon\leftarrow$ ExpandEllipse$(\varepsilon,\mathbf{p}_{*})$
15       $H\leftarrow$ ConstructHalfSpace$(\varepsilon,\mathbf{p}_{*})$
16       $\bm{O}_{b,\text{remain}}\leftarrow$ RemoveOrClipObstacles$(\bm{O}_{b,\text{remain}},H)$
@@ III-B2 Calculation Details
Once $\hat{\mathbf{p}}_{*}$ is obtained, the original closest point can be recovered as $\mathbf{p}_{*}=\mathcal{F}(\hat{\mathbf{p}}_{*})$. The optimization problem (16) is convex, as both $\mathcal{F}^{-1}(O)$ and $C^{-1}(B)$ are convex sets, and it can be solved using general convex optimization techniques. However, solving it this way may be computationally expensive. To improve efficiency, we leverage the GJK algorithm, which iteratively determines the support (extreme) point in a given direction, updates the search direction, and refines the simplex to approximate the closest point, ensuring fast convergence. For simplicity, the support function to search the support point for any convex set $\bullet$ in a direction $\mathbf{v}$ is defined as:
@@ 9      $\varsigma\leftarrow\{w\}\cup\varsigma$
10       $\hat{\mathbf{p}}_{*},\varsigma\leftarrow$ FindClosestPoint$(\varsigma)$
11       $\varsigma\leftarrow$ UpdateSimplex$(\varsigma)$
@@ 13             $\hat{\mathbf{p}}_{*}\leftarrow$ origin
In the ellipse generation step, once the closest point in inverse space $\hat{\mathbf{p}}_{*}$ is found, we can quickly verify whether the original point $\mathbf{p}_{*}=\mathcal{F}(\hat{\mathbf{p}}_{*})$ lies inside the original ellipse by checking $\|\hat{\mathbf{p}}_{*}\|_{2}<1$. The minor axis is then updated by substituting $\mathbf{p}_{*}$ into the ellipse equation while keeping the major axis fixed.

### v2

@@ II-A Vehicle Geometry and Dynamics
Figure 1: Vehicle geometry and state. $P$: rear-axle center; $O^{\prime}$: instantaneous center; FLFL: front-left corner; FRFR: front-right corner; RLRL: rear-left corner; RRRR: rear-right corner; $L_{f}$: distance from rear axle to front bumper; $L_{r}$: distance from rear axle to rear bumper; $W$: vehicle width.
@@ II-D3 Switching Points
where $i$ ranges from 1 to M1M-1.
@@ II-D4 Reference Trajectory Proximity
[…]
@@ II-D5 Collision Avoidance
A safe convex corridor is constructed around the reference trajectory using convex polygons. Let $P_{k}^{(i)}$ be the polygon enclosing $\mathbf{x}_{r,k}^{(i)}$, expressed as $P_{k}^{(i)}=\big{\{}\mathbf{p}\mid A_{k}^{(i)}\mathbf{p}\leq b_{k}^{(i)}\big{\}}$, where $A_{k}^{(i)}\in\mathbb{R}^{n\times 2}$ and $b_{k}^{(i)}\in\mathbb{R}^{n}$ are known matrices, and n{n} is the number of enclosing hyperplanes at this point. The method to construct the safe convex corridor will be presented in Section III-B. Collision avoidance is enforced by constraining all vehicle corners to remain within the polygon:
$A_{k}^{(i)}\mathbf{p}_{\bullet,k}^{(i)}\leq b_{k}^{(i)}+\sigma_{k}^{(i)}\cdot\mathbf{1},\sigma_{k}^{(i)}\geq 0,\text{for}\;\bullet\in\big{\{}\text{FL, FR, RL, RR}\big{\}}$ (10)
@@ 2 Set dynamical feasibility tolerance $e_{f}$
[…]
@@ 11 $P\leftarrow\mathbb{R}^{2}$
[…]
@@ III-B2 Calculation Details
[…]
@@ 9 $\varsigma\leftarrow\{w\}\cup\varsigma$
[…]
@@ 13    $\hat{\mathbf{p}}_{*}\leftarrow$ origin
[…]